Water of mass mw=1kg in a calorimeter has an initial temperature of Tw=200K. A hot piece of metal of mass mx=0.5kg and temperature of Tx=600K is placed in the calorimeter and the system is allowed to go to equilibrium with a final temperature of Tf=500K. What is the specific heat of the metal, cx? Use cw=4000J/kgK.
Answer #1

Heat lost by metal = heat gained by water

0.5 * cx * ( 600 - 500) = 1 * 4000 * ( 500 - 200)

0.5 * cx * 100 = 1200,000

cx = 24000 J / kg.K